OpenStudy (anonymous):

Simplify. Express the product as a radical expression.. (x^3)^(3/4)

OpenStudy (anonymous):

is it

OpenStudy (cgreenwade2000):

Break down the exponent (3/4) into 3*(1/4) ex x^3/4 \[=\sqrt[4]{x ^{3}}\]

OpenStudy (anonymous):

but thats not one of the options

OpenStudy (cgreenwade2000):

I didn't answer the problem for you. I just showed an example.

OpenStudy (anonymous):

Ooobut am i right?

OpenStudy (cgreenwade2000):

No.

OpenStudy (anonymous):

Yy not?

OpenStudy (cgreenwade2000):

\[x ^{2}\times x ^{2/5}= x ^{2} \sqrt[5]{x ^{2}}\]

OpenStudy (cgreenwade2000):

That's another example.

OpenStudy (anonymous):

the last one?

OpenStudy (cgreenwade2000):

An exponent as a fraction means the denominator is the root and the numerator just stays as a regular exponent

OpenStudy (cgreenwade2000):

I'm not going to tell you which one is right.

OpenStudy (anonymous):

ok thanks